OpenStudy (yacoub1993):

What is the simplest form of the expression? ∛45 + ∛2058 - ∛750 A. 9∛6 B. 2.8∛6 C. 4∛6 D. 14∛6

OpenStudy (yacoub1993):

A

OpenStudy (shubhamsrg):

are you sure first term if 45 ? Please check and confirm

OpenStudy (anonymous):

\[\sqrt[3]{45} + (7 + 5) \sqrt[3]{6} \]

OpenStudy (yacoub1993):

yes http://prntscr.com/34tsfh

OpenStudy (anonymous):

Check first term like @shubhamsrg said

OpenStudy (yacoub1993):

there is the question captured

OpenStudy (anonymous):

\[1.96\sqrt[3]{6} + 2\sqrt[3]{6} \]

OpenStudy (anonymous):

\[\approx 4\sqrt[3]{6} \]